# Tradeoffs Between Alignment and Helpfulness in Language Models with Representation Engineering

### v4

@@ A.2 Discussion of assumptions
Linear classification with margin $\Delta$ (assumption 1): We expect the representation engineered vectors $r_{e}$ to be good classifiers because they are obtained by methods of finding directions in the latent space that maximize the distance between representations of positive and negative textual statements. For example, in Zou et al. (2023a) the first principle component is used as a steering vector, obtained via $pca_{1}=argmax_{v}\{\mathbb{E}_{good,bad}[|\langle v,r_{good}-r_{bad}\rangle|^{2}]\}$ and in Jorgensen et al. (2023) the steering vector is obtained as the average of difference between positive and negative statements $\frac{1}{N}\sum_{i=1}^{N}(r^{i}_{good}-r^{i}_{bad})$. In these examples, $r_{good}$ and $r_{bad}$ are representations of queries and not the latent space embedding of the answers, as in the definition of $\Delta$-representation-separability, but we expect the steering vectors to behave similarly on them. In subsection A.3, we show that indeed $\delta r_{e}(q)$ clusters positive and negative responses to harmful queries in the model’s latent space. In appendix F we also formulate a theorem equivalent to theorem 1, but with an imperfect classifier.
Random directionality of last hidden layer change (assumption 2): When answering queries that are unrelated to the behavior being enhanced by representation engineering, the directionality of the injected vectors are expected to be random w.r.t. the representations of the answers to the query. Therefore, the highest probability tokens are expected to be injected with random noise. We validate this in the next subsection, by looking at the noise injected into the top 10 highest probability tokens in knowledge queries (which typically make over 90% of the probability mass).
@@ A.3 Experiments for assumptions
Here we empirically check the validity of our assumptions and empirically estimate the values of the parameters in the bounds. The experiments were performed on Llama 2 13B and Llama 2 13B chat. We first verify a linear relation between the representation engineering coefficient $r_{e}$ to the last hidden layer change of assumption 3, which yields $\lambda$. Then, we verify the normal distribution assumption 2 and the linear classification of assumption 1.
@@
In figures 8 and 9 we plot the change in norm for Llama 2 13B chat (injected with racist vectors) and Llama 2 13B (injected with not racist vectors) respectively, on the datasets “international law", “medical genetics" and “high school computer science". We add fitted curves to estimate $\lambda$. We find that it is in the range $40-60$.
@@
In figure 12 and 13, we plot the distance between the centers of representation clusters for positive and negative answers to harmful queries as the norm of harmful vectors is increased, for Llama 2 13B chat and Llama 2 13B respectively. As can be seen, the distance between the clusters increases, which corresponds to an increase in $\mathbb{E}[\langle\delta r_{e}(q),U^{T}(e_{good}-e_{bad})\rangle]$. We can define a range of coefficients in which the increase is bounded from below by a linear curve of the form in equation 11, meaning that the change in the model’s representation separates the positive and negative answer representations, similarly to the definition of $\Delta$-representation separability, but with mean instead of min. Thus by equation 11, the tangent of the lower bounding lines of figures 12 and 13 are an estimate for $\Delta\lambda$. From, this we get that $\Delta\lambda$ is approximately $0.1-0.3$
@@ Appendix B Proof of theorem 1
The theorem utilizes assumptions 3 and 1. The behavior expectation is:
@@
Where $r(q)$ is the final hidden layer representation and $\delta r(q)$ is the change to the last hidden layer due to representation engineering on the previous layers. $a_{+}\in good$ and $a_{-}\in bad$ denote the aligned and misaligned answers respectively, i.e.  $B(a_{\pm})=\pm 1$.
@@
$<\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)\cdot\min_{a^{\prime}_{-}\in bad}\{exp(\langle\delta r(q),U^{T}e_{a^{\prime}_{-}}\rangle)\}}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}\rangle)\cdot\max_{a^{\prime}_{+}\in good}exp(\langle\delta r(q),U^{T}e_{a^{\prime}_{+}}\rangle)}=$ (17)
$=\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}\rangle)}\cdot\frac{1}{\max_{a^{\prime}_{+}\in good,a_{-}\in bad}exp(\langle\delta r(q),U^{T}e_{a^{\prime}_{+}}\rangle)}$ (18)
$=\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}-U^{T}e_{a_{-}}\rangle)}\cdot\frac{1}{exp(\max_{a^{\prime}_{+}\in good,a_{-}\in bad}\langle\frac{\delta r(q)}{|\delta r(q)|},U^{T}e_{a^{\prime}_{+}}\rangle\cdot|\delta r(q)|)}$ (19)
From $\Delta$ margin linear classification of $\{U^{T}a_{+}\}_{a_{+}\in good}$ and $\{U^{T}a_{-}\}_{a_{-}\in good}$ by $\frac{\delta r(q)}{|\delta r(q)|}$ (assumption 1):
@@ Appendix C Proof of theorem 2
The theorem utilizes assumptions 3 and 2. Notice that:
@@
Denote $X_{i}=\langle\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|},U^{T}e_{i}\rangle$ and by $P_{correct}^{0}$ the probability of answering correctly without representation engineering:
@@
Denote by $I_{\pm}=\{i\in incorrect(\epsilon)|\pm X_{i}-X_{correct}>0\}$. Using the AM-GM inequality:
$\leq\frac{P^{0}_{correct}}{P^{0}_{correct}+\sum_{i\in I_{+}}P_{i}^{0}e^{\frac{\sum_{i\in I_{+}}P_{i}^{0}(X_{i}-X_{correct})}{\sum_{i\in I_{+}}P_{i}^{0}}|\delta r_{e}(q)|}+\sum_{i\in I_{-}}P_{i}^{0}e^{\frac{\sum_{i\in I_{-}}P_{i}^{0}(X_{i}-X_{correct})}{\sum_{i\in I_{-}}P_{i}^{0}}|\delta r_{e}(q)|}}$ (32)
@@
$\leq\frac{P^{0}_{correct}}{P^{0}_{correct}+\min\{P_{-},P_{+}\}(1+\frac{1}{2}\min\{c_{-},c_{+}\}^{2}|\delta r_{e}(q)|^{2})}$ (35)
@@
$\leq\frac{P^{0}_{correct}}{P^{0}_{correct}+\min\{P_{-},P_{+}\}(1+\frac{1}{2}\min\{c_{-},c_{+}\}^{2}\lambda^{2}|r_{e}|^{2})}$ (36)
Under the assumption that $X_{i}$ distribute randomly (assumption 2), $c_{\pm}$ are a weighted sum of positive/negative random variables with parameter $\sigma$, which we can refactor to $\sigma^{2}\cdot c_{\pm}^{\prime 2}$ where $c_{\pm}^{\prime}$ are the same but with $\sigma^{\prime}=1$. Yielding:
@@
Notice that with probability $\frac{1}{T}$ the set $I_{\pm}$ is empty, therefor with probability $1-\frac{2}{T}$ both sets are not empty, thus $P_{\pm}>0$ and $c_{+}>0$, $c_{-}<0$.
Notice that $P(X_{correct}>X_{i})=\frac{1}{2}$, thus $i\in I_{+}$ with probability $\frac{1}{2}$. Therefor $P_{+}$ is a weighted sum of Bernoulli variables with weights $\{P^{0}_{i}\}_{i\in incorrect}$.
Appendix D Proof of Corollary 2
@@
Appendix E Experimental Details
E.1 Compute
@@
E.2 Obtaining the engineered representations
@@
  • For the harmless vectors for Llama 2 13B, we take harmful instructions from AdvBench dataset, create pairs of harmless and harmful responses to them, sampled from Llama2 13B chat. The created pairs of prompts follow the format: “[INST]Question:[INSTRUCTION] Answer: [RESPONSE(Llama2-13b-chat)][/INST]". The responses of Llama2 13b chat were sampled by injecting the model with harmfulness vectors to produce the misaligned answer, and the aligned answers were sampled without representation engineering. For Llama 3.1 8B, the usual method used for harmful vectors sufficed, only this time the desired behavior is opposite, so we subtract the representations of the prompts in reverse order.
@@ E.4 Helpfulness experiments
While the bound of theorem 2 is with probability $1-\frac{2}{|V|}=\frac{1}{2}$ in the case of $4$ answers, as explained in D, for the other $\frac{2}{|V|}$ probability, the helpfulness is bounded with equal probability either by a sigmoid or by a reverse sigmoid, such that together they contribute approximately $\frac{1}{|V|}$ to the expectation value of the helpfulness (due to their small overlap), leading to corollary 2, in which the average helpfulness converges to $\frac{1}{|V|}=\frac{1}{4}$ in the case of our experiment, as can be seen in figure 5. Around $r_{e}=0$, the contribution of these sigmoids to the helpfulness expectation value can be bounded with the parabolic bound of theorem 2 as shown in the proof provided in appendix D. Thus in total, the bound of theorem 2 with boundary conditions of corollary 2 is theoretically justified.
@@ E.5 Figures
Appendix F Relaxation to Soft Margin
In the proof of theorem 1, we use the assumption that the change to the last hidden layer representation due to representation engineering linearly classifies the representations of positive and negative answers to a query with margin $\Delta$ (as explained in appendix A). We can relax this assumption by assuming that some of the negative (positive) responses’ representations, are misclassified as aligned (misaligned) answers by $\delta r_{e}(q)$, in the sense that:
@@ Appendix F Relaxation to Soft Margin
Where $B_{0}=B[P_{\theta}(\cdot|q)]$ is the behavior expectation without representation engineering and $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.
Proof.
@@
Appendix G Relation of Representation engineering to Finetuning with Preference Learning
To a degree one can draw a relation between representation engineering and preference learning.
@@ Appendix G Relation of Representation engineering to Finetuning with Preference Learning
For an LLM, one iteration of gradient descent on the preference learning loss with learning rate $\eta$ is equivalent to representation engineering with coefficient $r_{e}=\eta$.
Proof.
@@
Thus at each layer, the representation is shifted in a direction that maximizes the difference between positive and negative responses’ representations, $U^{T}(e_{y_{+}}-e_{y_{-}})$. Which is equivalent to representation engineering with coefficient $r_{e}=\eta$, and vectors $R_{e}=\{\mathbb{E}_{(x,y^{+},y^{-})\sim D}[\frac{\partial r(x)}{\partial r^{l}(x)}\cdot U^{T}(e_{y_{+}}-e_{y_{-}})]\}_{l=1}^{L}$
Appendix H Extension of Results Beyond Binary Behavior Score
@@
Where $B_{0}=B[P_{\theta}(\cdot|q)]$ and $P_{+}$ are the behavior expectation and probability of aligned answer without representation engineering, and $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.
The behavior bound has a different form, but it behaves the same – for $r_{e}=0$, it coincides with $B_{0}$, around $r_{e}=0$ it is linear, and for $r_{e}\rightarrow\infty$ it approaches $+1$. The proof, presented in H.1, essentially follows the proof of theorem 1, except besides the $P_{\pm}$ terms (probability mass of positive and negative responses without representation engineering) there is also a $P_{0}$ term.
@@
Where $P_{\pm}$ are the probabilities of aligned/misaligned answers without representation engineering, and $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.
Here we see that the behavior expectation converges to the maximal score $b_{+}$, for which $\delta r_{e}^{(L)}$ can classify answers below and above the score. The trend is similar to theorem 1, with a sigmoidal behavior, but without the tightness on behavior expectation at $r_{e}=0$, due to the more complex behavior scoring function. The proof is presented in H.2.
H.1 Proof of theorem 4
Following the same proof as in 1, up to equation 22, but replacing the sum over negative answers to sum over negative and neutral answers, we obtain by denoting $P_{\pm}$, the sum over positive/negative answers without representation engineering, and by $P_{0}$ sum over neutral answers:
@@ H.1 Proof of theorem 4
H.2 Proof of theorem 5
Following the same proof idea as in theorem 1, starting with equation 12 but replacing the scores in the numerator for positive and negative answers with $b_{+}$ and $-1$ (for worst case), up to equation 22, denote by $P_{+}$ the probability without representation engineering for answers with score $>b_{+}$ and by $P_{-}$ the rest:
@@ H.2 Proof of theorem 5
Appendix I Extension of Results to Multi-Token Answers
@@
I.1 Alignment
Starting with alignment, we note that if the model is limited to producing $N$ tokens, then from corollary 1, we can ensure that with a large enough representation engineering coefficient, each token will correspond to an aligned response:
@@ I.1 Alignment
We see that larger coefficients of representation engineering improve the behavior expectation, similarly to corollary 1, but with multiple token answers. By inverting the relation between $r_{e}$ and $\epsilon$, and placing it in the behavior expectation bound, we obtain a sigmoid-like behavior, that is linear for $r_{e}\approx 0$.
@@ I.2 Helpfulness
Let $P_{\theta}$ be a language model and $q$ be a query with answer $a=a_{1}...a_{N}$ containing at most $N$ tokens. Denote by $\{P_{0}^{i}\}_{i=1}^{N}$ the probability assigned to each correct token $\{a_{i}\}_{i=1}^{N}$ in the sequence without representation engineering, such that the probability of the full sequence is $P_{0}=\prod_{i=1}^{N}P_{0}^{i}$. Then under the conditions of theorem 2 holding at each decoding step, we have with probability of at least $1-\frac{2N}{T}$:

### v5

@@ A.2 Discussion of assumptions
Linear classification with margin $\Delta$ (assumption 2): We expect the representation engineered vectors $r_{e}$ to be good classifiers because they are obtained by methods of finding directions in the latent space that maximize the distance between representations of positive and negative textual statements. For example, in Zou et al. (2023a) the first principle component is used as a steering vector, obtained via $pca_{1}=argmax_{v}\{\mathbb{E}_{good,bad}[|\langle v,r_{good}-r_{bad}\rangle|^{2}]\}$ and in Jorgensen et al. (2023) the steering vector is obtained as the average of difference between positive and negative statements $\frac{1}{N}\sum_{i=1}^{N}(r^{i}_{good}-r^{i}_{bad})$. In these examples, $r_{good}$ and $r_{bad}$ are representations of queries and not the latent space embedding of the answers, as in the definition of $\Delta$-representation-separability, but we expect the steering vectors to behave similarly on them. In subsection A.3, we show that indeed $\delta r_{e}(q)$ clusters positive and negative responses to harmful queries in the model’s latent space. In appendix H we also formulate a theorem equivalent to theorem 1, but with an imperfect classifier.
Random directionality of last hidden layer change (assumption 3): When answering queries that are unrelated to the behavior being enhanced by steering, the directionality of the injected vectors are expected to be random w.r.t. the representations of the answers to the query. Therefore, the highest probability tokens are expected to be injected with random noise. We validate this in the next subsection, by looking at the noise injected into the top 10 highest probability tokens in knowledge queries (which typically make over 90% of the probability mass).
@@ A.3 Experiments for assumptions
Here we empirically check the validity of our assumptions and empirically estimate the values of the parameters in the bounds. The experiments were performed on Llama 2 13B and Llama 2 13B chat. We first verify a linear relation between the steering vector coefficient $r_{e}$ to the last hidden layer change of assumption 1, which yields $\lambda$. Then, we verify the normal distribution assumption 3 and the linear classification of assumption 2.
@@
In figures 7 and 8 we plot the change in norm for Llama 2 13B chat (injected with racist vectors) and Llama 2 13B (injected with not racist vectors) respectively, on the datasets “international law", “medical genetics" and “high school computer science". We add fitted curves to estimate $\lambda$. We find that it is in the range $40-60$.
@@
In figure 11 and 12, we plot the distance between the centers of representation clusters for positive and negative answers to harmful queries as the norm of harmful vectors is increased, for Llama 2 13B chat and Llama 2 13B respectively. As can be seen, the distance between the clusters increases, which corresponds to an increase in $\mathbb{E}[\langle\delta r_{e}(q),U^{T}(e_{good}-e_{bad})\rangle]$. We can define a range of coefficients in which the increase is bounded from below by a linear curve of the form in equation 11, meaning that the change in the model’s representation separates the positive and negative answer representations, similarly to the definition of $\Delta$-representation separability, but with mean instead of min. Thus by equation 11, the tangent of the lower bounding lines of figures 11 and 12 are an estimate for $\Delta\lambda$. From, this we get that $\Delta\lambda$ is approximately $0.1-0.3$. In section 4, we obtained values of $\Delta\lambda$ in the rage $0.5-3$ from the free parameter fit on the bound of theorem 2 to the data. The difference between these two ranges is attributed to the method of the empirical estimation of $\Delta$ from the linear classification condition that looks for an upper bound on it on the entire $r_{e}$ range, while the main change in alignment in figure 4 occurs in a more specific range, where the upper bound of $\Delta$ is evidently bigger.
@@ Appendix B Proof of theorem 1
The theorem utilizes assumptions 1 and 2. The behavior expectation is:
@@
Where $r(q)$ is the final hidden layer representation and $\delta r(q)$ is the change to the last hidden layer due to steering on the previous layers. $a_{+}\in good$ and $a_{-}\in bad$ denote the aligned and misaligned answers respectively, i.e.  $B(a_{\pm})=\pm 1$.
@@
$<\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)\cdot\max_{a^{\prime}_{-}\in bad}\{exp(\langle\delta r(q),U^{T}e_{a^{\prime}_{-}}\rangle)\}}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}\rangle)\cdot\min_{a^{\prime}_{+}\in good}exp(\langle\delta r(q),U^{T}e_{a^{\prime}_{+}}\rangle)}=$ (17)
Moving the maximum in the numerator to the denominator turns it into a minimum and the exponent’s argument becomes negative, we obtain a product of two minimum terms, which we can jointly write as:
$=\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}\rangle)}\cdot\frac{1}{\min_{a^{\prime}_{+}\in good,a_{-}\in bad}exp(\langle\delta r(q),U^{T}e_{a^{\prime}_{+}}-U^{T}e_{a^{\prime}_{-}}\rangle)}$ (18)
As the exponent is a monotonic function, we can insert the minimum into the exponent:
$=\frac{\sum_{a_{-}\in bad}exp(\langle r(q),U^{T}e_{a_{-}}\rangle)}{\sum_{a_{+}\in good}exp(\langle r(q),U^{T}e_{a_{+}}\rangle)}\cdot\frac{1}{exp(\min_{a^{\prime}_{+}\in good,a_{-}\in bad}\langle\frac{\delta r(q)}{|\delta r(q)|},U^{T}e_{a^{\prime}_{+}}-U^{T}e_{a^{\prime}_{-}}\rangle\cdot|\delta r(q)|)}$ (19)
From $\Delta$ margin linear classification of $\{U^{T}a_{+}\}_{a_{+}\in good}$ and $\{U^{T}a_{-}\}_{a_{-}\in good}$ by $\frac{\delta r(q)}{|\delta r(q)|}$ (assumption 2), the minimum in the denominator is larger than $\Delta$:
@@ Appendix C Proof of theorem 2
The theorem utilizes assumptions 1 and 3. Notice that:
@@
Denote $X_{i}=\langle\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|},U^{T}e_{i}\rangle$ and by $P_{correct}^{0}$ the probability of answering correctly without steering:
@@
Denote by $I_{\pm}=\{i\in incorrect(\epsilon)|\pm(X_{i}-X_{correct})>0\}$ (i.e. $X_{i}$’s that are larger/smaller than $X_{correct}$). Also denote by $P_{i}^{0}=P_{\theta}(a_{i}|q)$ and $Y_{i}=|\delta r_{e}(q)|(X_{i}-X_{correct})$. We obtain two sums of the form $\sum_{i\in I_{\pm}}P_{i}e^{Y_{i}}$. Since the exponent is a convex function, using Jensen’s inequality, on the sums yields $\sum_{i\in I}P_{i}e^{Y_{i}}\geq(\sum_{i\in I}P_{i})\cdot e^{\frac{\sum_{j\in I}P_{j}Y_{i}}{\sum_{j\in I}P_{j}}}$. Plugging this in:
$\leq\frac{P^{0}_{correct}}{P^{0}_{correct}+(\sum_{i\in I_{+}}P_{i}^{0})\cdot e^{\frac{\sum_{j\in I_{+}}P_{j}^{0}(X_{j}-X_{correct})}{\sum_{j\in I_{+}}P_{j}^{0}}|\delta r_{e}(q)|}+(\sum_{i\in I_{-}}P_{i}^{0})\cdot e^{\frac{\sum_{j\in I_{-}}P_{j}^{0}(X_{j}-X_{correct})}{\sum_{j\in I_{-}}P_{j}^{0}}|\delta r_{e}(q)|}}$ (32)
@@
$\leq\frac{P^{0}_{correct}}{P^{0}_{correct}+\min\{P_{-},P_{+}\}(1+\frac{1}{2}\min\{|c_{-}|,c_{+}\}^{2}|\delta r_{e}(q)|^{2})}$ (35)
@@
$\leq\frac{P^{0}_{correct}}{P^{0}_{correct}+\min\{P_{-},P_{+}\}(1+\frac{1}{2}\min\{|c_{-}|,c_{+}\}^{2}\lambda^{2}|r_{e}|^{2})}$ (36)
Under the assumption that $X_{i}$ distribute randomly (assumption 3), $c_{\pm}$ are a weighted sum of positive/negative random variables with parameter $\sigma$, which we can refactor to $\sigma\cdot c_{\pm}^{\prime}$ where $c_{\pm}^{\prime}$ are the same variables but normalized to $\sigma^{\prime}=1$. Denoting $\beta=\min\{|c^{\prime}_{-}|,c^{\prime}_{+}\}$, yields:
@@
Notice that $I_{-}$ is empty if $X_{i}>X_{correct}$ for all $i\in incorrect(\epsilon)$, and from assumption 3, these random variables are identically distributed, hence from symmetry, the event that $X_{correct}$ is the smallest of the $T$ random variables is $1/T$. Thus, with probability $\frac{1}{T}$ the set $I_{\pm}$ is empty, therefore with probability $1-\frac{2}{T}$ both sets are not empty, thus $P_{\pm}>0$ and $c_{+}>0$, $c_{-}<0$.
From the above-mentioned symmetry arising from the random variables $X_{correct},\{X_{i}\}_{i\in I_{-}\cup I_{+}}$ being identically distributed, for each individual $i$, $P(X_{correct}>X_{i})=\frac{1}{2}$, thus $i\in I_{+}$ with probability $\frac{1}{2}$. Therefore, $P_{+}$ is a weighted sum of Bernoulli variables with weights $\{P^{0}_{i}\}_{i\in incorrect}$.
Appendix D Alignment Guarantee with Steering
In contrast to Wolf et al. (2023), that has a framework centralized on using prompts to misalign frozen models, i.e. models whose weights and representations are not changed after training, here the model is not frozen due to steering, and accordingly a different result is obtained on guaranteeing an aligned response – for any adversarial attack, using large enough norms with representation engineering produces an aligned response if the learned injected representations accumulate to a good classifier of positive and negative answer representations in the final layer. This is formalized here as a corollary of theorem 1.
Corollary 1.
Let $\epsilon>0$, $P_{\theta}$ a language model and $q$ a prompt that induces negative behavior $B[P_{\theta}(\cdot|q)]<\gamma<0$ without steering. Under the conditions of theorem 1, using an injected vector norm of $r_{e}>\frac{1}{\Delta\lambda}(arctanh(1-\epsilon)-arctanh(\gamma))$ leads to behavior expectation $B[P_{\theta,r_{e}}(\cdot|q)]>1-\epsilon$.
Appendix E Helpfulness at the Limit of Large Steering Vectors
When considering the average helpfulness over a dataset in a scenario where the number of answers is constant, $N$ (such as multiple choice questions), we obtain that on average, the model will converge to answering $1/N$ of the questions correctly as steering is increased:
Corollary 2.
proof:
@@ Appendix E Helpfulness at the Limit of Large Steering Vectors
Thus at the limit of $r_{e}\rightarrow\infty$, we obtain a bound of $1/T$.
Appendix F Experimental Details
F.1 Compute
@@
F.2 Obtaining the engineered representations
@@
  • For the harmless vectors for Llama 2 13B and Llama 3.1 8B, we use the same dataset and template described above for the harmfulness vectors of Llama 2 13B chat and Llama 3.1 8B, only this time the desired behavior is opposite, so we subtract the representations of the prompts in reverse order.
@@ F.4 Helpfulness experiments
While the bound of theorem 2 is with probability $1-\frac{2}{|V|}=\frac{1}{2}$ in the case of $4$ answers, as explained in E, for the other $\frac{2}{|V|}$ probability, the helpfulness is bounded with equal probability either by a sigmoid or by a reverse sigmoid, such that together they contribute approximately $\frac{1}{|V|}$ to the expectation value of the helpfulness (due to their small overlap), leading to corollary 2, in which the average helpfulness converges to $\frac{1}{|V|}=\frac{1}{4}$ in the case of our experiment, as can be seen in figure 5. Around $r_{e}=0$, the contribution of these sigmoids to the helpfulness expectation value can be bounded with the parabolic bound of theorem 2 as shown in the proof provided in appendix E. Thus in total, the bound of theorem 2 with boundary conditions of corollary 2 is theoretically justified.
@@ F.5 Figures
Appendix G Helpfulness Experiments on Code Generation
In section 4, we showed the model’s helpfulness on knowledge based question answering as a function of steering satisfies theorem 2. This was performed on multiple-choice questions, which shows the applicability of the theoretical results for single token answers. For demonstrating the theoretical results on tasks requiring generation of full sequences, we test the model’s coding skills with the humaneval dataset (Chen et al., 2021). As can be seen in figure 17, The model’s performance is peaked around $r_{e}=0$, and it decays parabolically ar $r_{e}$ increases, as predicted in theorem 2. We note that the asymmetry between positive and negative coefficients is captured in our theoretical bounds.
Appendix H Relaxation to Soft Margin
In the proof of theorem 1, we use the assumption that the change to the last hidden layer representation due to steering linearly classifies the representations of positive and negative answers to a query with margin $\Delta$ (as explained in appendix A). We can relax this assumption by assuming that some of the negative (positive) responses’ representations, are misclassified as aligned (misaligned) answers by $\delta r_{e}(q)$, in the sense that:
@@ Appendix H Relaxation to Soft Margin
Where $B_{0}=B[P_{\theta}(\cdot|q)]$ is the behavior expectation without steering and $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.
Proof:
@@
Appendix I Relation of Steering to Finetuning with Preference Learning
To a degree one can draw a relation between steering and preference learning.
@@ Appendix I Relation of Steering to Finetuning with Preference Learning
For an LLM, one iteration of gradient descent on the preference learning loss with learning rate $\eta$ is equivalent to steering with coefficient $r_{e}=\eta$.
Proof:
@@
Thus at each layer, the representation is shifted in a direction that maximizes the difference between positive and negative responses’ representations, $U^{T}(e_{y_{+}}-e_{y_{-}})$. Which is equivalent to steering with coefficient $r_{e}=\eta$, and vectors $R_{e}=\{\mathbb{E}_{(x,y^{+},y^{-})\sim D}[\frac{\partial r(x)}{\partial r^{l}(x)}\cdot U^{T}(e_{y_{+}}-e_{y_{-}})]\}_{l=1}^{L}$
Appendix J Extension of Results Beyond Binary Behavior Score
@@
Where $B_{0}=B[P_{\theta}(\cdot|q)]$ and $P_{+}$ are the behavior expectation and probability of aligned answer without steering, and $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.
The behavior bound has a different form, but it behaves the same – for $r_{e}=0$, it coincides with $B_{0}$, around $r_{e}=0$ it is linear, and for $r_{e}\rightarrow\infty$ it approaches $+1$. The proof, presented in J.1, essentially follows the proof of theorem 1, except besides the $P_{\pm}$ terms (probability mass of positive and negative responses without steering) there is also a $P_{0}$ term.
@@
Where $P_{\pm}$ are the probabilities of aligned/misaligned answers without steering, and $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.
Here we see that the behavior expectation converges to the maximal score $b_{+}$, for which $\delta r_{e}^{(L)}$ can classify answers below and above the score. The trend is similar to theorem 1, with a sigmoidal behavior, but without the tightness on behavior expectation at $r_{e}=0$, due to the more complex behavior scoring function. The proof is presented in J.2.
J.1 Proof of theorem 4
Following the same proof as in 1, up to equation 22, but replacing the sum over negative answers to sum over negative and neutral answers, we obtain by denoting $P_{\pm}$, the sum over positive/negative answers without steering, and by $P_{0}$ sum over neutral answers:
@@
J.2 Proof of theorem 5
Following the same proof idea as in theorem 1, starting with equation 12 but replacing the scores in the numerator for positive and negative answers with $b_{+}$ and $-1$ (for worst case), up to equation 22, denote by $P_{+}$ the probability without steering for answers with score $>b_{+}$ and by $P_{-}$ the rest:
@@
Appendix K Extension of Results to Multi-Token Answers
@@
K.1 Alignment
Starting with alignment, we note that if the model is limited to producing $N$ tokens, then from corollary 1, we can ensure that with a large enough steering coefficient, each token will correspond to an aligned response:
@@
We see that larger coefficients of steering improve the behavior expectation, similarly to corollary 1, but with multiple token answers. By inverting the relation between $r_{e}$ and $\epsilon$, and placing it in the behavior expectation bound, we obtain a sigmoid-like behavior, that is linear for $r_{e}\approx 0$.
@@
Let $P_{\theta}$ be a language model and $q$ be a query with answer $a=a_{1}...a_{N}$ containing at most $N$ tokens. Denote by $\{P_{0}^{i}\}_{i=1}^{N}$ the probability assigned to each correct token $\{a_{i}\}_{i=1}^{N}$ in the sequence without steering, such that the probability of the full sequence is $P_{0}=\prod_{i=1}^{N}P_{0}^{i}$. Then under the conditions of theorem 2 holding at each decoding step, we have with probability of at least $1-\frac{2N}{T}$: